# From Source Properties to Strong-Field Tests: a multipronged analysis of GW250114 with an effective one-body model for generic orbits

### v1

@@ Abstract
As previously mentioned, this analysis, though similar in spirit to the pSEOBNR test performed by the LVK (Brito:2018rfr; Pompili:2025cdc; LIGOScientific:2025obp), differs in its implementation. pSEOBNR constrains deviations from the IMR-expected QNM frequencies, while we directly constrain deviations in the IMR-expected remnant properties that inform the QNM frequencies that is used for modelling the post-peak waveform. Therefore, to compare the results, we use the LVK’s publicly available pSEOBNR posterior samples and reconstruct the corresponding remnant properties. Starting from the intrinsic binary parameters, we use NR fitting formulae to obtain the GR-consistent inspiral predictions for the remnant, $\bigl(M_{f}^{\rm IMR}(1+z),\,\chi_{f}^{\rm IMR}\bigr)$. Independently, we map the measured quasi-normal–mode quantities $\bigl(f_{220}(1+\delta f_{220}),\,\tau_{220}(1+\delta\tau_{220})\bigr)$ to the corresponding remnant mass and spin, $\bigl(M_{f}^{\rm dev}(1+z),\,\chi_{f}^{\rm dev}\bigr)$, using the relations of Berti:2006wq. We then quantify departures from GR through $\delta X_{f}\equiv X_{f}^{\rm dev}/X_{f}^{\rm IMR}-1,$ for $X_{f}\in\{M_{f},\chi_{f}\}$. As shown in Fig. 6, the resulting pSEOBNR constraints include the GR prediction also within the $50\%$ credible interval. Our bounds are somewhat tighter, consistent with the fact that our analysis leverages information from the subdominant $(4,\pm 4)$ multipole in addition to the dominant mode.
@@
Our second test allows a fractional deviation in the peak amplitude of the dominant $(2,2)$ mode, $\delta A_{22}^{\rm peak}$, sampling from a uniform prior on (1,4)(-1,4) under the quasi-circular assumption. For such an assumption, symmetry imposes the relation $h_{\ell,-m}=(-1)^{\ell}h_{\ell,m}^{\ast}$, ensuring that the (2,2)(2,-2) mode is fully determined by the $(2,2)$ mode and removing additional freedom in the waveform morphology. We find that $\delta A_{22}^{\rm peak}=0$ lies within the $\sim$82% credible interval, consistent with GR (See Fig. 8) and has a posterior width of $\mathcal{O}(10^{-1})$. For comparison, this value is just one to two orders of magnitude larger than the typical systematic uncertainty in NR simulations obtained from different codes (see e.g. Fig. 8 of Rashti:2024yoc). This fact is rather remarkable, showcasing how GW observations are soon to enter the high precision regime.
@@
Discussion: Our complementary analysis confirms and strengthens LIGOScientific:2025epi; LIGOScientific:2025obp results, indicating GW250114 as a nearly mass-symmetric quasi-spherical BBH merger with total mass $M_{T}\simeq 66\,M_{\odot}$ and low spins $\chi_{1,2}\simeq 0.1$, decisively disfavouring eccentric configurations ($\mathcal{B}\sim 20$). These properties point to a first-generation binary formed through isolated field evolution (Postnov:2014tza) with $\geq 96\%$ probability. Additionally owing to these properties and the signal’s high SNR ($\gtrsim 75$), we obtain decisive evidence for the subdominant $(4,\pm 4)$ multipoles, detectable up to $\Delta t^{\mathrm{geo}}_{\mathrm{start}}=5M$ beyond the signal’s peak, thereby strengthening the finding of LIGOScientific:2025obp.
@@
We then subject GR to a trio of stringent, complementary tests in the strong-field, non-linear regime with this signal. First, we perform a residual analysis, subtracting the full ensemble of posterior-supported waveforms from the data, and confirm that TEOBResumS-Dalí fully captures the signal, with no statistically significant deviations from noise ($p\simeq 0.3$). Second, we carry out the one of the first direct parameterised test of the subdominant $(4,\pm 4)$ multipole’s amplitude consistency with GR  in a BBH signal, introducing a fractional amplitude deviation $\delta h_{44}$ and recovering its posterior from the data. We find that the GR value lies within the 90% credible region, but the $\delta h_{44}$ posterior exhibits a shifted bimodality, a structure reproduced in simulated signal analysis where a NRSur7dq4 waveform is recovered using TEOBResumS-Dalí, but not in matched $TEOBResumS-Dalí$ recoveries. The agreement between the signal and mismatched simulated signals points to waveform systematics in the $(4,\pm 4)$ mode of TEOBResumS-Dalí. Lastly, using the first parameterised eccentric BBH waveform model, we find a GR-consistent plunge–merger–ringdown portion, with the GR expectation $(\delta M_{f}=\delta\chi_{f}=0)$ lying well within the 50% credible region for both non-eccentric cases and $\delta A^{\mathrm{peak}}_{22}$ lying inside the 90% credible region. Regarding the latter, we note that while finalising the paper, we became aware of Grimaldi2025 that also constrains the GW250114 merger’s amplitude.
@@
Collectively these analyses, together with those in LIGOScientific:2025epi; LIGOScientific:2025obp, provide compelling and multifaceted evidence for the validity of GR in the dynamical, strong-field gravity of a BBH coalescence. GW250114 thus stands as a watershed event, heralding a new era of precision strong-gravity physics where future high-SNR detections will enable even more stringent constraints on potential deviations from GR.
@@
We analyse 8 s of publicly available Advanced LIGO data around the event (LIGOScientific:2025snk), sampled at 4096 Hz, using Bilby with the dynesty sampler and the TEOBResumS-Dalí waveform model in four different configurations (Ashton:2018jfp; Speagle:2019ivv): (1) non-eccentric, aligned-spin; (2) non-eccentric, precessing-spin; (3) eccentric, aligned-spin; and (4) eccentric, precessing-spin BBH mergers. This model-selection step identifies the most probable configuration a priori, thereby maximising the constraining power of subsequent deviation measurements  [^6] By picking the most probable configuration, we shrink the allowed parameter space, which makes the subsequent measurements give tighter constraints..
Following Romero-Shaw:2020owr, we adopt isotropic priors for sky position and orientation, and a distance prior uniform in comoving volume and source time. Configuration-specific priors are used for intrinsic parameters–particularly spins, eccentricity, and mean anomaly–evaluated at a reference frequency of 13.3 Hz. To probe deviations from GR, we use settings similar to those above, introducing uniform priors on the relevant deviation parameters as needed.
For our full-signal analyses, we use the Whittle likelihood, with the publicly available data power spectrum estimated on-source via a Bayesian approach that combines splines for broadband features and Lorentzians for narrow spectral lines (Cornish:2014kda). For measuring the persistence of the subdominant $(4,\pm 4)$ multipole in the post-inspiral signal, we instead employ a time-domain likelihood implemented in Bilby (Ashton:2018jfp); further details are provided below.
@@
Bayesian inference libraries like Bilby model the evenly sampled time-domain detector data $\bm{d}$ as a linear superposition of a possible GW signal $\bm{s}(\bm{\theta})$ and noise $\bm{n}$ (Ashton:2018jfp). Their goal is to identify a template $\bm{h}(\bm{\theta}\mid M)$ from a chosen waveform model $M$ (in our case, either TEOBResumS-Dalí or pTEOBResumS-Dalí) that best represents the signal, such that the residual $\bm{r}=\bm{d}-\bm{h}(\bm{\theta}\mid M)$ follows the noise’s statistical properties.
@@
Under these assumptions, the single-detector log-likelihood ratio between the signal and noise assertion for the full analysis is (Finn:1992wt; Veitch:2009hd)
@@
Equivalently, we may express the likelihood in the discrete time domain using the noise covariance matrix $C_{ik}$ (Isi:2021iql). In our analysis, we compute $C$ from the autocorrelation function obtained by inverse Fourier transforming the PSD $S_{j}$. Assuming wide-sense stationarity, $C$ is Toeplitz and positive definite; we perform a Cholesky decomposition,
@@
which we use to evaluate the log-likelihood ratio defined in Eq. (4) to probe the temporal structure of $(4,\pm 4)$ modes (Chandra:2025ipu).
Instead of performing a new stochastic sampling for this restricted window, we recompute the likelihood using Eq. (4) by resampling from the existing IMR posteriors (Rover:2006ni; Ashton:2025xba). This means the samples from the IMR analysis act as draws from a proposal distribution, while the restricted time-domain likelihood defines a new target distribution. The ratio of these two, expressed as a weight factor for each sample, reweights both the IMR posterior and the evidence, providing an approximation to the restricted-domain posterior without rerunning the full sampler.
@@
To assess the robustness of our GR baseline model, we compute mismatches between TEOBResumS-Dalí and NRSur7dq4 (Varma:2019csw), a quasi-spherical multipolar model built by interpolating NR simulations and widely regarded as the most NR-faithful waveform currently available. We use the sky-maximized detection statistics of Harry:2017weg, maximized over time and reference azimuth, as well as over an additional in-plane rotation angle Pratten:2020ceb. We assume a lower frequency cutoff of $20$ Hz and an upper cutoff of $900$ Hz, employing the Advanced LIGO design sensitivity curve (LIGOScientific:2014pky). We find mismatches of $\bar{\mathcal{F}}\sim 10^{-4}-10^{-3}$ ($10^{-3}-10^{-2}$) between the two models when considering only the $\ell=2$ (including up to $\ell=4$) modes for systems drawn from the posterior of GW250114 (see left panel of Fig. 9). These numbers translate into critical SNR thresholds of $\rho_{\mathrm{crit}}\sim 70-100$ (203020-30) below which the two models should be indistinguishable Damour:2010zb. Therefore, while the $\ell=2$ modes of the two models are mostly indistinguishable for GW250114, the inclusion of higher-order modes may in principle lead to measurable differences given the signal’s high SNR.
@@
To demonstrate that GW250114 is less likely to have formed through a hierarchical merger process, we conduct a simple Monte Carlo study informed by the GWTC–3 population estimates (KAGRA:2021duu). We generate an initial BBH population by drawing component masses $m_{k}$ from the Power Law + Peak model, and spin magnitudes $\chi_{k}$ and tilt angles $\theta_{k}$ from the Default spin model (Talbot:2017yur; Talbot:2018cva). We label the BHs in this sample as “1G”, while noting that the observed GWTC–3 population may itself contain hierarchically assembled remnants; thus, this population does not represent a pure 1G population aka BHs formed directly via stellar-collapse.
We merge a fraction of this 1G+1G population to form second-generation (2G) BHs. Their final masses and spins are computed using NR-informed fits (Hofmann:2016yih; Jimenez-Forteza:2016oae), and their recoil speeds are obtained using fits from Gerosa:2016sys. For subsequent steps, we retain only those remnants whose recoil speeds are below $500\,\mathrm{km\,s^{-1}}$, a threshold achievable only in a few host environments (e.g., AGN discs (Gerosa:2017kvu)); thus, our results provide optimistic upper bounds for hierarchical growth.

### v2

@@ Abstract
As previously mentioned, this analysis, though similar in spirit to the pSEOBNR test performed by the LVK (Brito et al., 2018; Pompili et al., 2025; Abac and others, 2025a), differs in its implementation. pSEOBNR constrains deviations from the IMR-expected QNM frequencies, while we directly constrain deviations in the IMR-expected remnant properties that inform the QNM frequencies that is used for modelling the post-peak waveform. Therefore, to compare the results, we use the LVK’s publicly available pSEOBNR posterior samples and reconstruct the corresponding remnant properties. Starting from the intrinsic binary parameters, we use NR fitting formulae to obtain the GR-consistent inspiral predictions for the remnant, $\bigl(M_{f}^{\rm IMR}(1+z),\,\chi_{f}^{\rm IMR}\bigr)$. Independently, we map the measured quasi-normal–mode quantities $\bigl(f_{220}(1+\delta f_{220}),\,\tau_{220}(1+\delta\tau_{220})\bigr)$ to the corresponding remnant mass and spin, $\bigl(M_{f}^{\rm dev}(1+z),\,\chi_{f}^{\rm dev}\bigr)$, using the relations of Berti and Cardoso (2006). We then quantify departures from GR through $\delta X_{f}\equiv X_{f}^{\rm dev}/X_{f}^{\rm IMR}-1,$ for $X_{f}\in\{M_{f},\chi_{f}\}$. As shown in Fig. 6, the resulting pSEOBNR constraints include the GR prediction also within the $50\%$ credible interval. Our bounds are somewhat tighter, consistent with the fact that our analysis leverages information from the subdominant $(4,\pm 4)$ multipole in addition to the dominant mode.
@@
Our second test allows a fractional deviation in the peak amplitude of the dominant $(2,2)$ mode, $\delta A_{22}^{\rm peak}$, sampling from a uniform prior on (1,4)(-1,4) under the quasi-circular assumption. For such an assumption, symmetry imposes the relation $h_{\ell,-m}=(-1)^{\ell}h_{\ell,m}^{\ast}$, ensuring that the (2,2)(2,-2) mode is fully determined by the $(2,2)$ mode and removing additional freedom in the waveform morphology. We find that $\delta A_{22}^{\rm peak}=0$ lies within the $\sim$82% credible interval, consistent with GR (See Fig. 8) and has a posterior width of $\mathcal{O}(10^{-1})$. For comparison, this value is just one to two orders of magnitude larger than the typical systematic uncertainty in NR simulations obtained from different codes (see e.g. Fig. 8 of Rashti et al. (2025)). This fact is rather remarkable, showcasing how GW observations are soon to enter the high precision regime.
@@
Discussion: Our complementary analysis confirms and strengthens Abac and others (2025c, a) results, indicating GW250114 as a nearly mass-symmetric quasi-spherical BBH merger with total mass $M_{T}\simeq 66\,M_{\odot}$ and low spins $\chi_{1,2}\simeq 0.1$, decisively disfavouring eccentric configurations ($\mathcal{B}\sim 20$). These properties point to a first-generation binary formed through isolated field evolution (Postnov and Yungelson, 2014) with $\geq 96\%$ probability. Additionally owing to these properties and the signal’s high SNR ($\gtrsim 75$), we obtain decisive evidence for the subdominant $(4,\pm 4)$ multipoles, detectable up to $\Delta t^{\mathrm{geo}}_{\mathrm{start}}=5M$ beyond the signal’s peak, thereby strengthening the finding of Abac and others (2025a).
@@
We then subject GR to a trio of stringent, complementary tests in the strong-field, non-linear regime with this signal. First, we perform a residual analysis, subtracting the full ensemble of posterior-supported waveforms from the data, and confirm that TEOBResumS-Dalí fully captures the signal, with no statistically significant deviations from noise ($p\simeq 0.3$). Second, we carry out the one of the first direct parameterised test of the subdominant $(4,\pm 4)$ multipole’s amplitude consistency with GR  in a BBH signal, introducing a fractional mode deviation and recovering its posterior from the data. We find that the GR value lies within the 90% credible region, but the deviation posterior exhibits a mild shift in amplitude, a structure reproduced in simulated signal analysis where a NRSur7dq4 waveform is recovered using TEOBResumS-Dalí, but not in matched $TEOBResumS-Dalí$ recoveries. The agreement between the signal and mismatched simulated signals points to waveform systematics in the $(4,\pm 4)$ mode of TEOBResumS-Dalí. Lastly, using the first parameterised eccentric BBH waveform model, we find a GR-consistent plunge–merger–ringdown portion, with the GR expectation $(\delta M_{f}=\delta\chi_{f}=0)$ lying well within the 50% credible region for both non-eccentric cases and $\delta A^{\mathrm{peak}}_{22}$ lying inside the 90% credible region. Regarding the latter, we note that while finalising the paper, we became aware of Grimaldi and others (2025) that also constrains the GW250114 merger amplitude.
@@
Collectively these analyses, together with those in Abac and others (2025c, a), provide compelling and multifaceted evidence for the validity of GR in the dynamical, strong-field gravity of a BBH coalescence. GW250114 thus stands as a watershed event, heralding a new era of precision strong-gravity physics where future high-SNR detections will enable even more stringent constraints on potential deviations from GR.
@@
We analyse 8 s of publicly available Advanced LIGO data around the event (Abac and others, 2025d), sampled at 4096 Hz, using Bilby with the dynesty sampler and the TEOBResumS-Dalí waveform model in four different configurations (Ashton and others, 2019; Speagle, 2020): (1) non-eccentric, aligned-spin; (2) non-eccentric, precessing-spin; (3) eccentric, aligned-spin; and (4) eccentric, precessing-spin BBH mergers. This model-selection step identifies the most probable configuration a priori, thereby maximising the constraining power of subsequent deviation measurements  [^6] By picking the most probable configuration, we shrink the allowed parameter space, which makes the subsequent measurements give tighter constraints..
Following Romero-Shaw and others (2020), we adopt isotropic priors for sky position and orientation, and a distance prior uniform in comoving volume and source time. Configuration-specific priors are used for intrinsic parameters–particularly spins, eccentricity, and mean anomaly–evaluated at a reference frequency of 13.3 Hz. To probe deviations from GR, we use settings similar to those above, introducing uniform priors on the relevant deviation parameters as needed.
For our full-signal analyses, we use the Whittle likelihood, with the publicly available data power spectrum estimated on-source via a Bayesian approach that combines splines for broadband features and Lorentzians for narrow spectral lines (Cornish and Littenberg, 2015). For measuring the persistence of the subdominant $(4,\pm 4)$ multipole in the post-inspiral signal, we instead employ a time-domain likelihood implemented in Bilby (Ashton and others, 2019); further details are provided below.
@@
Bayesian inference libraries like Bilby model the evenly sampled time-domain detector data $\bm{d}$ as a linear superposition of a possible GW signal $\bm{s}(\bm{\theta})$ and noise $\bm{n}$ (Ashton and others, 2019). Their goal is to identify a template $\bm{h}(\bm{\theta}\mid M)$ from a chosen waveform model $M$ (in our case, either TEOBResumS-Dalí or pTEOBResumS-Dalí) that best represents the signal, such that the residual $\bm{r}=\bm{d}-\bm{h}(\bm{\theta}\mid M)$ follows the noise’s statistical properties.
@@
Under these assumptions, the single-detector log-likelihood ratio between the signal and noise assertion for the full analysis is (Finn, 1992; Veitch and Vecchio, 2010)
@@
Equivalently, we may express the likelihood in the discrete time domain using the noise covariance matrix $C_{ik}$ (Isi and Farr, 2021). In our analysis, we compute $C$ from the autocorrelation function obtained by inverse Fourier transforming the PSD $S_{j}$. Assuming wide-sense stationarity, $C$ is Toeplitz and positive definite; we perform a Cholesky decomposition,
@@
which we use to evaluate the log-likelihood ratio defined in Eq. (4) to probe the temporal structure of $(4,\pm 4)$ modes (Chandra and Calderón Bustillo, 2025).
Instead of performing a new stochastic sampling for this restricted window, we recompute the likelihood using Eq. (4) by resampling from the existing IMR posteriors (Rover et al., 2006; Ashton, 2025). This means the samples from the IMR analysis act as draws from a proposal distribution, while the restricted time-domain likelihood defines a new target distribution. The ratio of these two, expressed as a weight factor for each sample, reweights both the IMR posterior and the evidence, providing an approximation to the restricted-domain posterior without rerunning the full sampler.
@@
To assess the robustness of our GR baseline model, we compute mismatches between TEOBResumS-Dalí and NRSur7dq4 (Varma et al., 2019), a quasi-spherical multipolar model built by interpolating NR simulations and widely regarded as the most NR-faithful waveform currently available. We use the sky-maximized detection statistics of Harry et al. (2018), maximized over time and reference azimuth, as well as over an additional in-plane rotation angle Pratten and others (2021). We assume a lower frequency cutoff of $20$ Hz and an upper cutoff of $900$ Hz, employing the Advanced LIGO design sensitivity curve (Aasi and others, 2015). We find mismatches of $\bar{\mathcal{F}}\sim 10^{-4}-10^{-3}$ ($10^{-3}-10^{-2}$) between the two models when considering only the $\ell=2$ (including up to $\ell=4$) modes for systems drawn from the posterior of GW250114 (see left panel of Fig. 9). These numbers translate into critical SNR thresholds of $\rho_{\mathrm{crit}}\sim 70-100$ (203020-30) below which the two models should be indistinguishable Damour et al. (2011). Therefore, while the $\ell=2$ modes of the two models are mostly indistinguishable for GW250114, the inclusion of higher-order modes may in principle lead to measurable differences given the signal’s high SNR.
@@
To demonstrate that GW250114 is less likely to have formed through a hierarchical merger process, we conduct a simple Monte Carlo study informed by the GWTC–3 population estimates (Abbott and others, 2023). We generate an initial BBH population by drawing component masses $m_{k}$ from the Power Law + Peak model, and spin magnitudes $\chi_{k}$ and tilt angles $\theta_{k}$ from the Default spin model (Talbot and Thrane, 2017, 2018). We label the BHs in this sample as “1G”, while noting that the observed GWTC–3 population may itself contain hierarchically assembled remnants; thus, this population does not represent a pure 1G population aka BHs formed directly via stellar-collapse.
We merge a fraction of this 1G+1G population to form second-generation (2G) BHs. Their final masses and spins are computed using NR-informed fits (Hofmann et al., 2016; Jiménez-Forteza et al., 2017), and their recoil speeds are obtained using fits from Gerosa and Kesden (2016). For subsequent steps, we retain only those remnants whose recoil speeds are below $500\,\mathrm{km\,s^{-1}}$, a threshold achievable only in a few host environments (e.g., AGN discs (Gerosa and Berti, 2017)); thus, our results provide optimistic upper bounds for hierarchical growth.